OpenStudy (anonymous):

The measurement of one angle of a right triangle is 61°. What is the measurement of the missing angle?

OpenStudy (anonymous):

Right triangle has one angle that is 90 degrees. The other angle is 61 degrees. So, 90+61=151 Now, assuming you know that all of a triangle's angle's adds up to 180 degrees, you would then proceed to take 180 degrees and subtract that by 151. 180-151=29 29 degrees would be the final angle and your answer. To check your work add the three angles to see if they add up to 180: 90+61+29=180